# Towards Robust and Realistic Human Pose Estimation via WiFi Signals

### v2

@@ 3.2 Domain-Consistent Representation Learning
$\mathcal{L}_{\mathrm{Mask}}=||\mathbf{X}^{\prime}_{i}-\mathbf{\hat{X}}_{i}||_{2}^{2}.$ (1)
@@ 4.1 Dataset
Table 1: State-of-the-art comparisons on MM-Fi dataset regarding 3D pose estimation. The best and the second-best results are marked in Red and Blue, respectively.
@@ 4.3 Ablation Study
Influence of Pre-training Components. Table 5 studies the influence of each component in our pre-training phase. Notably, our per-training phase yields more general and robust representations than the model trained from scratch. Moreover, the temporal-consistent contrastive strategy contributes a lot, highlighting the necessity of capturing motion dynamics in WiFi-based pose estimation.
@@
${}^{\phantom{\dagger}}$ ${}^{\phantom{\dagger}}$
${}^{\phantom{\dagger}}$ ${}^{\phantom{\dagger}}$
${}^{\phantom{\dagger}}$ ${}^{\phantom{\dagger}}$
@@ A.2 Evaluation Metric
Percentage of Correct Keypoints (PCK@$\alpha$ (%)):; Measure the percentage of predictions that fall within a certain threshold distance from the ground truth. The threshold is set as a fraction $\alpha$ of the torso length following the previous works [10, 45]. It is widely used to evaluate the local accuracy.
@@ A.4 Joints Analysis
Table 10: State-of-the-art comparisons on MM-Fi dataset regarding 2D pose estimation under Protocol 3 with Setting 1 split. The best and the second-best results are marked in Red and Blue, respectively.

### v3

@@ 3.2 Domain-Consistent Representation Learning
[…]
@@ 4.1 Dataset
Table 1: State-of-the-art comparisons on $M$-Fi dataset regarding 3D pose estimation. The best and the second-best results are marked in Red and Blue, respectively.
@@ 4.3 Ablation Study
Influence of Pre-training Components. Table 4.3 studies the influence of each component in our pre-training phase. Notably, our per-training phase yields more general and robust representations than the model trained from scratch. Moreover, the temporal-consistent contrastive strategy contributes a lot, highlighting the necessity of capturing motion dynamics in WiFi-based pose estimation.
@@
[…]
@@ A.2 Evaluation Metric
[…]
@@ A.4 Joints Analysis
Table 10: State-of-the-art comparisons on $M$-Fi dataset regarding 2D pose estimation under Protocol 3 with Setting 1 split. The best and the second-best results are marked in Red and Blue, respectively.